40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 249     Entry time: Fri Jan 18 15:31:47 2008
Author: rob 
Type: Update 
Category: LSC 
Subject: Thursday's locking 

rob, johnnie, andrey

On Thursday night we got the intereferometer fully locked in a power-recycled FPMI state. The obstacles included the REFL166 phase being wrong by 180 deg (because that's the correct phase for DRMI locking) and getting confused (again) by the "manual" mode dewhite switching at the ETMs. After turning on the dewhites and the MICH correction, we took the noise spectrum below.
Attachment 1: DARMnoise080118.png  50 kB  | Hide | Hide all
DARMnoise080118.png
ELOG V3.1.3-